# Mixture of Gaussian-distributed Prototypes with Generative Modelling for Interpretable and Trustworthy Image Recognition

### v2

@@ 8       Update memory queue $\mathcal{B}^{c}$ for each class $c$;
III-F MGProto Pruning by Prototype Importance Prior
@@
In this work, we propose to remove less-important or trivial prototypes in a trained MGProto model, resorting to the inherent mixture constraint of GMMs: $\sum_{m=1}^{M}\boldsymbol{\pi}_{m}^{c}=1$ for each class $c$, as mentioned in Section III-B. To be specific, we take advantage of the prototype importance prior $\boldsymbol{\pi}_{m}^{c}$, where a large (or low) prior means the respective prototype is important (or trivial) to characterise the underlying class-conditional distribution $p(\mathbf{x}|c)$. Based on the above analysis, we prune a trained MGProto by keeping only important prototypes corresponding to the top-$\widetilde{M}$ priors for each class, where $\widetilde{M}<M$. This can avoid choosing complicated per-class pruning thresholds, which is required in previous pruning schemes [10, 16].
IV Experiments
IV-A Datasets and Evaluation Metrics
@@
IV-B Implementation Details
We perform experiments on various CNN backbone architectures: VGG16 (V16), VGG19 (V19), ResNet34 (R34), ResNet50 (R50), ResNet152 (R152), DenseNet121 (D121), and DenseNet161 (D161), which are all pre-trained on ImageNet [73], except for ResNet50 on CUB that is pre-trained on iNaturalist [74]. In accordance with prior studies, the prototype dimension $D=64$ and prototype number $M=10$ for all backbones and datasets. Following [25, 17], we obtain more fine-grained feature maps with $\bar{H}=\bar{W}=14$ by dropping the final max-pooling layer in the backbones. Then two add-on $1\times 1$ convolutional layers (without activation function) are appended to reduce the number of feature channels match the prototype dimension. The memory capacity $N$ is set to 800, 1000, 2000, 2000 for CUB, Cars, Dogs, and Pets, respectively. In Eq. (12), we have $\lambda_{1}$ = 0.2, and $\lambda_{2}$ is set as 0.5 according to [39]. For the prototype mining, we use $T$ = 20 levels of active patches.
@@ IV-B Implementation Details
IV-C Comparison With SOTA Methods
IV-C1 Accuracy of Interpretable Image Recognition
@@
IV-C2 Accuracy of Trustworthy Image Recognition
@@
Figure 8: Interpretable and trustworthy reasoning of our MGProto method for (a) classifying an In-Distribution Le Conte Sparrow image from CUB test set and (b) identifying an Out-of-Distribution Keeshond image from Pets. We only show the top-2 classification predictions.
IV-C3 Pruning of MGProto Model
Table V shows the MGProto pruning results. We first attempt to apply the purity-based pruning strategy [10] to our MGProto, but only a few prototypes are removed. This means that the strategy is improper to our method, mostly because our prototypes have large purity, evidenced in Table VII. On the other hand, if the merge-based pruning [41] is adopted, the model’s performance will be heavily sacrificed despite a substantial number of pruned prototypes.
@@ IV-C3 Pruning of MGProto Model
TABLE V: Model pruning results of our MGProto on CUB-200-2011, using different pruning strategies. % Pruned denotes the ratio between the number of pruned prototypes to that of the initial ones.
@@
IV-C4 Interpretability Quantification
Table VI shows the interpretability results evaluated on CUB. MGProto has the highest consistency score, indicating that the object parts activated by our prototypes are consistent across different images. This is likely because our prototypes are holistically learned from a weighted average of a large set of relevant features, allowing them to consistently capture cross-image semantics. MGProto also achieves competitive stability score. The lowest OIRR score demonstrates that MGProto relies more on the object region and less on the background context to support its classification decision. The DAUC result indicates that MGProto produces interpretations that best influence its classification predictions. We also compute the prototype purity in Table VII, showing that the prototypes produced by our MGProto have a high degree of purity, i.e., each prototype focuses on semantically-consistent object parts among different images.
TABLE VI: Quantitative interpretability results (%) on CUB test set using full images, where all models are based on VGG19 backbone.
@@ IV-C4 Interpretability Quantification
TABLE VII: Prototype purity (%) on train and test sets of CUB using full images, where all models are based on ResNet50 backbone.
@@
IV-D Ablation Studies
IV-D1 Auxiliary Loss and Prototype Mining
In Table VIII, we provide extensive ablation experiments on the auxiliary loss and prototype mining. Firstly, with the use of auxiliary loss $\mathcal{L}_{aux}$, the classification accuracy is improved, affirming its ability to aid the model backbone in enhancing feature extraction, as in [39]. However, $\mathcal{L}_{aux}$ does not have much effect on the interpretability measures, because it is designed primarily for feature extraction rather than interpretability.
@@ IV-D1 Auxiliary Loss and Prototype Mining
To explore the generality of our horse-racing based prototype mining strategy, we also apply it to other existing prototypical-part methods, e.g., ProtoPNet, TesNet, and ST-ProtoPNet. Results in Table IX show that our prototype mining strategy is still effective in these existing models relying on point-based learning techniques. We notice the accuracy improvement is particularly significant in the vanilla ProtoPNet [10], mostly because this baseline model has the worse classification performance among the three models.
TABLE VIII: Ablation analysis of the compared mining loss $\mathcal{L}_{mn}$ in Eq. (10) and auxiliary loss $\mathcal{L}_{aux}$ in Eq. (13) for the proposed MGProto method.
@@
TABLE IX: Improved accuracy (%) of existing prototypical-part networks on CUB, integrated with our prototype mining, based on ResNet34 backbone.
@@
IV-D2 Prototype Diversity in M-step
In Eq. (8), we introduce a diversity constraint to improve the within-class distances between prototype means. Table X presents an ablation study using full images of CUB and Cars. Our approach in Eq. (8) can significantly increase the average pair-wise distances ($L_{2}$) between prototype means. Also, prototypes with greater diversity tend to correlate with higher classification accuracy, in line with findings observed in established methods [15, 25, 45].
TABLE X: Effect of the prototype diversity constraint in Eq. (8) on the within-class distance between prototype means and classification accuracy (%).
@@ IV-D2 Prototype Diversity in M-step
IV-D3 Effect of Memory Bank
@@
IV-D4 $T$ and $\lambda_{1}$ in Prototype Mining
We also study the the effect of the two important hyperparameters $T$ and $\lambda_{1}$ utilised in our prototype mining, with results shown in Fig. 11. It can be seen that when $T$ is greater than 20, the accuracy of MGProto is high and generally stable. For the mining loss weight $\lambda_{1}$, setting it too low fails to introduce sufficient mining supervision signals, whereas increasing it too much also causes decreased classification accuracy. Therefore, we have $T=20$ and $\lambda_{1}=0.2$ in all other experiments.
@@ IV-D4 $T$ and $\lambda_{1}$ in Prototype Mining
V Conclusion
This work presented MGProto, a generative learning paradigm to produce mixtures of Gaussian-distributed prototypes for interpretable and trustworthy image recognition. Different from current point-based prototype learning models, MGProto naturally mitigates the performance degradation caused by the projection of point-based learned prototypes, and effectively recognises OoD inputs to ensure decision trustworthiness. Additionally, a novel and generic strategy is proposed to enhance the prototype learning by mining from less-salient object parts, inspired by the classic legend of Tian Ji’s horse-racing. MGProto can be substantially compressed via an importance-based pruning strategy, while still keeping high classification accuracy.
One limitation of the proposed method is that we empirically use the same number of prototypes (Gaussian components in GMM) for each class. In practice, the prototype number should vary across classes, given the different representation difficulties and imbalanced training samples among classes. In the future, we will explore more flexible techniques to learn class-adaptive and even class-agnostic prototypes for our method. Also, considering that our prototypes are learned to model the underlying data distributions, it is possible to sample from the distributions and generate new examples for counterfactual explanations. We plan to also explore this in our future work.

### v3

@@ 8       Update memory queue $\mathcal{B}^{c}$ for each class $c$;
3.6 MGProto Pruning by Prototype Importance Prior
@@
In this work, we propose to remove less-important or trivial prototypes in a trained MGProto model, resorting to the inherent mixture constraint of GMMs: $\sum_{m=1}^{M}\boldsymbol{\pi}_{m}^{c}=1$ for each class $c$, as mentioned in Section 3.2. To be specific, we take advantage of the prototype importance prior $\boldsymbol{\pi}_{m}^{c}$, where a large (or low) prior means the respective prototype is important (or trivial) to characterise the underlying class-conditional distribution $p(\mathbf{x}|c)$. Based on the above analysis, we prune a trained MGProto by keeping only important prototypes corresponding to the top-$\widetilde{M}$ priors for each class, where $\widetilde{M}<M$. This can avoid choosing complicated per-class pruning thresholds, which is required in previous pruning schemes [10, 16].
4 Experiments
4.1 Datasets and Evaluation Metrics
@@
4.2 Implementation Details
We perform experiments on various CNN backbone architectures: VGG16 (V16), VGG19 (V19), ResNet34 (R34), ResNet50 (R50), ResNet152 (R152), DenseNet121 (D121), and DenseNet161 (D161), which are all pre-trained on ImageNet [76], except for ResNet50 on CUB that is pre-trained on iNaturalist [77]. In accordance with prior studies, the prototype dimension $D=64$ and prototype number $M=10$ for all backbones and datasets. Following [25, 17], we obtain more fine-grained feature maps with $\bar{H}=\bar{W}=14$ by dropping the final max-pooling layer in the backbones. Then two add-on $1\times 1$ convolutional layers (without activation function555 Two convolutional layers without a non-linear activation between them are functionally similar to applying a single convolutional layer. Here we choose the two-layer structure for consistency with many prior ProtoPNet-based methods. ) are appended to reduce the number of feature channels to match the prototype dimension, thereby improving the computational efficiency. The memory capacity $N$ is set to 800, 1000, 2000, 2000 for CUB, Cars, Dogs, and Pets, respectively. In Eq. (12), we have $\lambda_{1}$ = 0.2, and $\lambda_{2}$ is set as 0.5 according to [40]. For the prototype mining, we use $T$ = 20 levels of active patches.
@@ 4.2 Implementation Details
4.3 Comparison With SOTA Methods
4.3.1 Accuracy of Interpretable Image Recognition
@@
4.3.2 Accuracy of Trustworthy Image Recognition
@@
Notice that the OoD detection is a long-standing research topic, and mainstream approaches harness post-hoc analysis, which builds a scoring criterion (e.g., logit-based [82, 64], distance-based [83, 84], and density-based [85, 86]) to indicate the ID-ness of the input. Relying on these scores, subsequent studies boost the OoD detection with auxiliary training data from outlier exposure [87] or synthesis [88]. To explore where the ProtoPNet-based methods stand in comparison with specialised OoD detection approaches, we also experiment on the standard CIFAR-10 benchmark. Consistent with the setup in [84, 89], we utilise CIFAR-10 [90] as ID dataset and the same model backbone (ResNet-18) in our method. OoD datasets include: SVHN [91], Places365 [92], LSUN [93], iSUN [94], and Textures [95]. In line with prior post-hoc approaches, we employ only ID data for model’s training, without using any auxiliary OoD training data. Experimental results in Table V demonstrate that our distribution-based prototypes achieve the best results on Textures dataset and deliver comparable average performances on par with other leading OoD detection approaches. In particular, our MGProto method greatly outperforms GEM [85] that models the class-conditional density using only a single Gaussian distribution, since our method leverages mixture of Gaussian distributions, enabling a more accurate characterization of the true underlying data density for each class.
TABLE V: Comparison with specialised OoD detection approaches, evaluated with FPR95 (%) and AUROC (%). We show the results on each individual OoD dataset and the results on average across all five OoD datasets.
Method OoD Dataset
SVHN Places365 LSUN iSUN Textures Average
FPR95↓ AUROC↑ FPR95↓ AUROC↑ FPR95↓ AUROC↑ FPR95↓ AUROC↑ FPR95↓ AUROC↑ FPR95↓ AUROC↑
MSP [96] 59.66 91.25 62.46 88.64 51.93 92.73 54.57 92.12 66.45 88.50 59.01 90.65
ODIN [97] 20.93 95.55 63.04 86.57 31.92 94.82 33.17 94.65 56.40 86.21 41.09 91.56
Energy [64] 54.41 91.22 42.77 91.02 23.45 96.14 27.52 95.59 55.23 89.37 40.68 92.67
Mahalanobis [83] 9.24 97.80 83.50 69.56 67.73 73.61 6.02 98.63 23.21 92.91 37.94 86.50
ViM [98] 24.95 95.36 63.04 86.57 7.26 98.53 33.17 94.65 56.40 86.21 36.96 92.26
ReAct [99] 48.16 92.32 37.25 93.13 18.09 96.91 20.35 95.59 96.51 47.41 34.25 94.09
KNN [84] 24.53 95.96 50.90 89.14 25.29 95.69 25.55 95.26 27.57 94.71 30.77 94.15
ConjNorm [86] 18.71 96.48 53.44 89.18 22.20 95.95 22.64 95.87 25.80 95.11 28.56 94.52
ASH [100] 28.94 94.84 27.29 91.31 9.06 98.34 21.61 95.95 35.02 93.63 27.29 94.81
GEM [85] 37.41 93.22 38.40 91.50 10.81 97.90 12.36 97.57 28.88 94.38 25.57 94.91
GEN [101] 26.10 93.83 37.40 93.42 8.99 98.39 24.09 95.03 26.62 94.13 24.64 94.96
NCA [89] 14.33 96.05 26.73 91.85 19.28 96.83 29.88 94.86 17.03 95.64 21.45 95.05
MGProto 28.56 94.38 39.40 91.38 10.74 98.03 28.20 94.75 12.26 97.59 23.83 95.23
@@
Figure 8: Interpretable and trustworthy reasoning of our MGProto method for (a) classifying an In-Distribution Le Conte Sparrow image from CUB test set and (b) identifying an Out-of-Distribution Keeshond image from Pets. In the figure, we only show the top-2 classification predictions.
4.3.3 Pruning of MGProto Model
Table VI shows the MGProto pruning results. We first attempt to apply the purity-based pruning strategy [10] to our MGProto, but only a few prototypes are removed. This means that the strategy is improper to our method, mostly because our prototypes have large purity, evidenced in Table VIII. On the other hand, if the merge-based pruning [42] is adopted, the model’s performance will be heavily sacrificed despite a substantial number of pruned prototypes.
@@ 4.3.3 Pruning of MGProto Model
TABLE VI: Model pruning results of our MGProto on CUB-200-2011, using different pruning strategies. % Pruned denotes the ratio between the number of pruned prototypes to that of the initial ones.
@@
4.3.4 Interpretability Quantification
Table VII shows the interpretability results evaluated on CUB. MGProto has the highest consistency score, indicating that the object parts activated by our prototypes are consistent across different images. This is likely because our prototypes are holistically learned from a weighted average of a large set of relevant features, allowing them to consistently capture cross-image semantics. MGProto also achieves competitive stability score. The lowest OIRR score demonstrates that MGProto relies more on the object region and less on the background context to support its classification decision. The DAUC result indicates that MGProto produces interpretations that best influence its classification predictions. We also compute the prototype purity in Table VIII, showing that the prototypes produced by our MGProto have a high degree of purity, i.e., each prototype focuses on semantically-consistent object parts among different images.
TABLE VII: Quantitative interpretability results (%) on CUB test set using full images, where all models are based on VGG19 backbone.
@@ 4.3.4 Interpretability Quantification
TABLE VIII: Prototype purity (%) on train and test sets of CUB using full images, where all models are based on ResNet50 backbone.
@@
4.4 Ablation Studies
4.4.1 Auxiliary Loss and Prototype Mining
In Table IX, we provide extensive ablation experiments on the auxiliary loss and prototype mining. Firstly, with the use of auxiliary loss $\mathcal{L}_{aux}$, the classification accuracy is improved, affirming its ability to aid the model backbone in enhancing feature extraction, as in [40]. However, $\mathcal{L}_{aux}$ does not have much effect on the interpretability measures, because it is designed primarily for feature extraction rather than interpretability.
@@ 4.4.1 Auxiliary Loss and Prototype Mining
To explore the generality of our horse-racing based prototype mining strategy, we also apply it to other existing prototypical-part methods, e.g., ProtoPNet, TesNet, and ST-ProtoPNet. Results in Table X show that our prototype mining strategy is still effective in these existing models relying on point-based learning techniques. We notice the accuracy improvement is particularly significant in the vanilla ProtoPNet [10], mostly because this baseline model has the worse classification performance among the three models.
TABLE IX: Ablation analysis of the compared mining loss $\mathcal{L}_{mn}$ in Eq. (10) and auxiliary loss $\mathcal{L}_{aux}$ in Eq. (13) for the proposed MGProto method.
@@
TABLE X: Improved accuracy (%) of existing prototypical-part networks on CUB, integrated with our prototype mining, based on ResNet34 backbone.
@@
4.4.2 Prototype Diversity in M-step
In Eq. (8), we introduce a diversity constraint to improve the within-class distances between prototype means. Table XI presents an ablation study using full images of CUB and Cars. Our approach in Eq. (8) can significantly increase the average pair-wise distances ($L_{2}$) between prototype means. Also, prototypes with greater diversity tend to correlate with higher classification accuracy, in line with findings observed in established methods [15, 25, 46].
TABLE XI: Effect of the prototype diversity constraint in Eq. (8) on the within-class distance between prototype means and classification accuracy (%).
@@ 4.4.2 Prototype Diversity in M-step
4.4.3 Effect of Memory Bank
@@
4.4.4 $T$ and $\lambda_{1}$ in Prototype Mining
We also investigate the the effect of the two important hyperparameters $T$ and $\lambda_{1}$ utilised in our prototype mining approach, with results shown in Fig. 11. It can be seen that when $T$ is greater than 20, the accuracy of MGProto is high and generally stable. For the mining loss weight $\lambda_{1}$, setting it too low fails to introduce sufficient mining supervision signals, whereas increasing it too much also causes decreased classification accuracy. Therefore, we have $T=20$ and $\lambda_{1}=0.2$ in all other experiments.
@@ 4.4.4 $T$ and $\lambda_{1}$ in Prototype Mining
5 Discussion and Limitation
As outlined in Eq. (1) and (14), the existing prototypical-part networks and our MGProto method require to ground prototypes in the human-understandable space of training images. This replacement step aims to find the closest feature patch for each prototype by going through all training images periodically at a regular interval during the training process, as shown in Fig. 2. However, this computationally-intensive step can hinder the scalability of these models to datasets with millions of training samples, such as ImageNet [76] that has a large number of classes. Furthermore, the large number of classes may challenge the learning process because prototypes from different classes may become closer to one another, leading to ambiguities in interpretation and an increased risk of misclassification. In this regard, developing more efficient replacement strategy (e.g. ball representation of prototypical concepts in ProtoConcept [45]) and more effective learning algorithm for prototypes is highly desirable in future research.
Our proposed MGProto method leverages class-specific prototypes, which are not shared among classes to ensure clear interpretability for each class. Nevertheless, this approach may face challenges in scenarios involving a large number of classes (e.g., 1000 classes in ImageNet) due to the following two reasons: 1) the number of prototypes scales linearly with the number of classes, making it intractable; and 2) as the number of classes grows, the prototypes will become highly similar to each other and the explanation maps they generate for different classes will start to be nearly identical, thereby significantly reducing interpretability. One potential solution is to incorporate the idea of class-shared prototypes [42, 43], which partially merges prototypes for those classes with similar features. By sharing prototypes between classes, we can address scalability (it becomes sublinear in the number of classes) and interpretability (we have many shared prototypes and a few specific prototypes that are more interpretable for each class). We argue that this adaptation can be straightforwardly implemented in our method, by making some GMM components shared across related classes.
Our MGProto method evaluates the model interpretability using objective metrics in Table VII and Table VIII. Another possibility to assess the interpretability relies on user studies [102, 45], which directly gauge how intuitive the explanations appear to human users. To enhance our evaluation, we plan to conduct user-based experiments for our proposed method in the future. Additionally, considering that our prototypes are learned to model the underlying data distributions, it is possible to sample from the distributions and generate new examples for counterfactual explanations. We plan to also explore this in future study.
6 Conclusion
This work presented MGProto, a generative learning paradigm to produce mixtures of Gaussian-distributed prototypes for interpretable and trustworthy image recognition. Different from current point-based prototype learning models, MGProto naturally mitigates the performance degradation caused by the projection of point-based learned prototypes, and effectively recognises OoD inputs to ensure decision trustworthiness. Additionally, a novel and generic strategy is proposed to enhance the prototype learning by mining from less-salient object parts, inspired by the classic legend of Tian Ji’s horse-racing. MGProto can be substantially compressed via an importance-based pruning strategy, while still keeping high classification accuracy. Experimentally, the proposed method obtains SOTA performances on interpretable image recognition, OoD detection, and interpretability quantification. A series of ablation studies are also performed to validate the significance of our key claims and effectiveness of the critical model designs.
Acknowledgements
This work was supported by funding from the Australian Government under the Medical Research Future Fund - Grant MRFAI000090 for the Transforming Breast Cancer Screening with Artificial Intelligence (BRAIx) Project. G. Carneiro acknowledges the support by the Engineering and Physical Sciences Research Council (EPSRC) through grant EP/Y018036/1. The authors would like to thank the editor and the anonymous reviewers for their valuable comments and suggestions.